OpenStudy (anonymous):

What is the exact value of the square root of the quantity 98 over 12.? Simplify if possible. Answer 7 the square root of 6 all over 6. the square root of 21 over 3. 7 the square root of 6 all over 2. 7 the square root of 2 all over 12.

OpenStudy (anonymous):

Do something like this \[ > \sqrt{ 98 \over 12 } = \sqrt{ 49 \over 6 } = { 7 \over \sqrt{ 6 }} \]

OpenStudy (shane_b):

And then rationalize the denominator to get a matching answer.

OpenStudy (anonymous):

What do you mean

OpenStudy (anonymous):

multiply both numerator and denominator with \sqrt{6}

OpenStudy (anonymous):

So 7 √6 42 √21 --- x --- = --- then simplify which will equal ------ √6 √6 36 3

OpenStudy (shane_b):

I don't think so :)\[\frac{7}{\sqrt{6}}*\frac{\sqrt{6}}{\sqrt{6}}=\frac{7\sqrt{6}}{6}\]